OpenStudy (anonymous):

solve the equation. 6(w-5)-3w=w+10

OpenStudy (anonymous):

distribute the 6, combine like terms, isolate w

OpenStudy (anonymous):

i need help getting past 6w-30-3w=w+10

OpenStudy (anonymous):

do i subract 3w from 6w and 3w?

OpenStudy (anonymous):

combine the ws on the left. remove the w on the right. remove the -30 on the left. then get w by itself by dividing

OpenStudy (anonymous):

so 6w and 3w are now 9w?

jimthompson5910 (jim_thompson5910):

No, 6w minus 3w is 3w because 6w - 3w = (6-3)w = 3w So again, 6w - 3w = 3w for all values of w

OpenStudy (anonymous):

i'm confused

jimthompson5910 (jim_thompson5910):

You have 6w-30-3w=w+10 which you can rewrite as 6w-3w-30=w+10 From here, you simplify 6w-3w to get 3w

jimthompson5910 (jim_thompson5910):

Does that make sense?

OpenStudy (anonymous):

so it's now 3w=w+10?

jimthompson5910 (jim_thompson5910):

No it should be 3w-30=w+10

OpenStudy (anonymous):

okay now what do we do?

jimthompson5910 (jim_thompson5910):

you want to isolate w, so you want to get all the w terms on one side

jimthompson5910 (jim_thompson5910):

so why not get w on the left side

jimthompson5910 (jim_thompson5910):

3w-30=w+10 3w-30+30=w+10+30 3w+0=w+40 3w=w+40 3w-w = 40 2w = 40 2w/2 = 40/2 w = 20

jimthompson5910 (jim_thompson5910):

Let me know if that makes sense or not.

OpenStudy (anonymous):

yes it does thanks

jimthompson5910 (jim_thompson5910):

you're welcome